The number of sales per week of an $80 pair of jeans is 25. The number of weekly sales goes

up to 45 when the price is reduced by 30%.

(a) Find the percentage change in demand.

(b) Find the average elasticity of demand.


1
Expert's answer
2021-03-08T09:22:28-0500

(a)

"\%Q=\dfrac{Q_2-Q_1}{Q_1}\times100\%=\dfrac{45-25}{25}\times100\%=80\%."

(b)

"E_d=\dfrac{\Delta \%Q}{\Delta\%P}=\dfrac{80\%}{-30\%}=-2.67."